# Safety-Driven Response Adaptive Randomisation: An Application in Non-inferiority Oncology Trials

### v1

@@
(June 10, 2025)
The majority of response-adaptive randomisation (RAR) designs in the literature use efficacy data to dynamically allocate patients. Their applicability in settings where the efficacy measure is observable with a random delay, such as overall survival, remains challenging. This paper introduces a RAR design referred to as SAFER (Safety-Aware Flexible Elastic Randomisation) design , which uses early-emerging safety data to inform treatment allocation decisions in oncology trials. However, the design is applicable to a range of settings where it may be desirable to favour the arm demonstrating a superior safety profile. This is particularly relevant in non-inferiority trials, which aim to demonstrate an experimental treatment is not inferior to the standard of care, while offering advantages in terms of safety and tolerability. Consequently, an unavoidable and well-established trade-off arises for such designs: to balance the goals of preserving inferential efficiency for the primary non-inferiority outcome while incorporating safety considerations into the randomisation process through RAR. Our method, defines a randomisation procedure which prioritises the assignment of patients to better-tolerated arms and adjusts the allocation proportion according to the observed association between safety and efficacy endpoints. We illustrate our procedure through a comprehensive simulation study, inspired by the CAPP-IT Phase III oncology trial. Our results demonstrate that SAFER preserves statistical power even when efficacy and safety endpoints are weakly associated and offers power gains when a strong positive association is present. Moreover, the approach enables a faster/slower adaptation when efficacy and safety endpoints are temporally aligned/misaligned, respectively.
Keywords Project Optimus  $\cdot$ Patient-driven  $\cdot$ Quality Of Life  $\cdot$ Efficacy  $\cdot$ Adverse events  $\cdot$ Dose reduction  $\cdot$ Drug discontinuation  $\cdot$ Pharmacokinetic  $\cdot$ Pharmacodynamic  $\cdot$ Exposure–response
@@ 1 Introduction
In contemporary drug development, optimal dosing has become a central focus, aiming for a precise, patient-centred approach in oncology. Initiatives like the FDA’s Project Optimus [1] underscore this shift, emphasizing the best balance between drug’s efficacy and safety to maximize patient benefit and minimise toxicity. Traditionally, early-phase clinical trials primarily prioritise safety, often focusing on major adverse events, with efficacy being largely exploratory. As development progresses, later phases are driven by efficacy decisions, and while comprehensive safety data are collected, this information is not typically formally integrated into design decisions, for example through an adaptive design. This leads to a critical challenge: despite the goal of improving the overall benefit-risk profile and enabling patients to stay on treatment longer, late-phase designs often do not formally use accumulating safety data to adapt trial conduct to be better aligned with those aims.
@@
The aforementioned issue becomes particularly relevant in the context of the modern oncology treatment landscape, which has been transformed by the advent of targeted therapies, monoclonal antibodies, and immune checkpoint inhibitors. These novel agents, designed to modulate specific molecular pathways, frequently demonstrate plateauing efficacy at doses well below the maximum tolerated dose (MTD) [2, 3]. Moreover, their administration typically spans prolonged treatment cycles and is often accompanied by chronic, low-grade adverse events (AEs), including fatigue, gastrointestinal disturbances, and dermatologic reactions. These AEs, although not classified as dose-limiting in the short windows of phase I trials, may significantly impair patient adherence, quality of life, and long-term tolerability in the later phases [4, 5]. A pertinent example to illustrate this is sotorasib (LUMAKRAS), a KRAS G12C inhibitor approved for non-small cell lung cancer [6]. For this treatment a 960 mg dose was selected for pivotal trials. Subsequent analyses showed that a 240 mg dose offered similar efficacy with a better safety profile, leading the FDA to mandate a post-marketing dose optimization study.
@@
The above rationale motivates our interest in the following crucial question: how can a late phase adaptive design help identify alternative dosing strategies—such as lower or fractionated doses, different schedules or regiments—deliver, with comparable or superior efficacy with fewer side effects? This is particularly relevant to address the potential impact on phase II and III studies from dosing decisions from early trials often adopted on immediate and severe safety data. This is especially the case for agents with delayed efficacy signals and cumulative toxicity profiles that influence long-term adherence and outcomes. In light of the growing emphasis on toxicity management in oncology, we anticipate an increasing number of studies, both in early and late phase of development, will focus on evaluating alternative dosing regimens with improved safety profiles [7].
@@
Our work aims to address the research question presented above by leveraging Response-Adaptive Randomisation (RAR). RAR refers to a class of statistical methodologies used in clinical trial design to dynamically adjust the probability of assigning patients to different treatment arms based on accumulating outcome data [8]. By increasingly favouring arms showing early superior performance, RAR can enhance both patient outcomes and the statistical efficiency of trials [9]. This adaptability is particularly advantageous when outcomes are available shortly after treatment initiation, allowing real-time modification of allocation proportions. However, in trials where primary efficacy outcomes are delayed – such as overall survival in oncology – RAR becomes more challenging due to the unavailability of timely, informative data during enrolment. Several strategies have been proposed to address this, including the use of early surrogate endpoints that correlate with long-term outcomes [10, 11, 12]. On the contrary, this paper introduces a RAR design that, alternatively, uses early-emerging safety data to inform treatment allocation decisions in oncology trials.
@@
However, a critical component of any safety-driven RAR is the inclusion on the possible interplay between safety and efficacy endpoints. While some drugs can maintain or improve efficacy after dose reductions, others could show reduced therapeutic benefit. From a statistical perspective, RAR designs may involve a trade-off with statistical power. In fact, while equal allocation does not necessarily maximise power in superiority trials [13], a balanced 1:1 allocation ratio is generally optimal in non-inferiority studies, where the alternative hypothesis assumes equal means (and variances) in the control and experimental arms. To this regards, Baldi Antognini et al. (2020) [14] introduced a compound optimal allocation strategy in survival endpoint trials to address the trade-off between "ethical" and inferential concerns in the choice of the randomisation proportion. Although their work is aligned with our rationale, it is limited to single-outcome efficacy trials since it does not capture the complexity of a joint safety-efficacy decisions.
@@
Therefore, we propose an innovative Safety-Aware Flexible Elastic Randomisation (SAFER) design. SAFER dynamically adjusts patient allocation proportions based on accumulating safety endpoint data, aiming to strike a balance between patient safety, maximization of clinical benefit, and the exploitation of the association between efficacy and safety outcomes. This adaptive strategy is particularly suited for late-phase clinical trials, where it enhances decision-making while preserving statistical efficiency, especially with respect to study power.
@@
The remainder of this paper is structured as follows. In Section 2, we define the key components of the SAFER design: (i) the efficacy endpoint, the associated power function, and sample size calculation; (ii) the definition of the safety endpoint; (iii) the target allocation proportion and the RAR procedure; and (iv) the integration of efficacy and safety data for the estimation of the allocation proportion. Section 3 introduces the CAPP-IT study as a motivating example and presents its re-design in a non-inferiority framework using the SAFER approach. Specifically, we outline the rationale and report the results from six distinct simulation scenarios, aimed at evaluating the performance of the proposed design under varying conditions. Section 4 discusses the simulation findings, explores broader implications of the SAFER design, and concludes with a summary of key contributions and directions for future research.
2 Components of the SAFER design
2.1 Efficacy Endpoint (Power and Sample Size calculation)
@@
Our proposed design is specifically tailored to accommodate for the typical primary endpoint used in Phase II/III oncology trials. Specifically, we consider a time-to-event measure, such as Progression-Free Survival (PFS). Consequently, both the sample size and the primary analysis of our design will be expressed throughout this paper in terms of this designated endpoint. The sample size will be determined accordingly to ensure adequate statistical power for that main primary analysis. We will now formally outline this rationale, including the underlying assumptions and notation.
@@
Let $j=E,C$ the experimental treatment and control arms, respectively. The patients’ primary efficacy endpoint (e) i.e., their survival time ($Y^{(e)}$ ) is assumed to be independent and identically distributed (i.i.d.) with probability density function (p.d.f) when treatment $j$ is assigned:
@@
where $\theta_{j}^{(e)-1}=\lambda_{j}^{(e)}\in\mathbb{R}^{+}$ denotes the event rate. Therefore, $\theta_{j}^{(e)}$ represents the mean of the treatment effect, and $\theta_{j}^{(e)2}$ represents the variance of the treatment effect. We will focus our design on the case in which the study will be testing for the non-inferiority of E with respect to C, the non-inferiority hypothesis for the mean difference becomes [15, 16]
@@
$W=\frac{\hat{\theta}_{C}^{(e)}-\hat{\theta}_{E}^{(e)}-m}{\sqrt{Var(\theta_{C}^{(e)}-\theta_{E}^{(e)})}}\sim\mathcal{N}(0,1),$ (3)
@@
$Var(\theta_{C}^{(e)}-\theta_{E}^{(e)})=\frac{1}{\sqrt{E}}\sqrt{\hat{\theta}_{C}^{(e)2}+\hat{\theta}_{E}^{(e)2}}$ (4)
We also incorporate an early stopping feature into our designs, recognizing that this feature, when combined with RAR, can offer advantages over RAR alone [17]. For simplicity, we have focused on only two analyses: an interim analysis (k=1) and the final analysis (k=2). Therefore, we will provide $W_{k=1}$ and $W_{k=2}$, estimated based on the first group available at the interim analysis and the complete sample at the end of follow-up, respectively.
We estimate the treatment effect through the Cox model [18] such that $W$ is approximated by $(\hat{\beta}_{E}-m)/SE(\hat{\beta}_{E})$ where $\hat{\beta}_{E}$ is the estimated difference between arms, i.e., $(\hat{\theta}_{C}^{(p)}-\hat{\theta}_{E}^{(p)})$ and $SE(\hat{\beta}_{E})$ is the square root of $Var(\hat{\beta}_{E})$. Although both the Cox model and exponential survival models are consistent estimators under exponentiality, in practice, the Cox model often converges more quickly and accurately in simulations. This is primarily due to its reliance on partial likelihood, fewer distributional assumptions, and better numerical conditioning in the presence of non-informative censoring and finite samples [19, 20].
Moreover, to estimate the minimum number of observed progression events $P$ needed for the study, we construct a $H_{0}$ rejection region as $R=\left\{-\infty,z_{1-\beta}\right\}$, with
@@
where $\alpha$ is the significance level, $\beta$ is the Type II Error rate, $HR_{H_{1}}$ is the hazard ratio ($\lambda_{E}/\lambda_{C}$) under the alternative hypothesis, and $\pi_{E}$ is the realized allocation proportion to the experimental arm (E). We will let $\pi_{E}$ vary from 0.5 to 0.8. Therefore, for each $\pi_{E}$ and associated rejection region, we can calculate the minimum number of events $P$ needed to test the non-inferiority on the primary endpoint between treatments, based on the following power function definition [21].
[…]
After the calculation of the minimum number of events $P$ needed to achieve a study power of at least 0.8, we adjust the total number of events needed due to the application of the Lan-DeMets O’Brien-Fleming $\alpha$-spending function [22] to account for $k=2$ analyses. Finally, due to the survival primary endpoint, the overall sample size ($N$) is estimated by dividing the total number of events needed ($P$) by the probability to observe the event during the follow-up, given the event rate expected in the control and experimental arms, under the alternative hypothesis.
@@ 2.2 Safety Endpoint
The safety endpoint in our design is the time to dose-reduction/drug discontinuation due to low drug tolerability or AEs. The patient’s safety endpoint (s) ($Y^{(s)}$), is assumed to be i.i.d., with p.d.f. when treatment $j$ is assigned:
@@
2.3 Allocation Proportion Target
For our proposed design, the Neyman allocation [23, 24, 25] is a reasonable choice of allocation proportion to target. This is because Neyman allocation under these assumptions will consistently assign the majority of patients to the treatment that patients are observed to discontinue later [14] (or tolerate for the longest). The Neyman allocation proportion (or the proportion to the treatment arm $\pi_{E}$) can be estimated based on estimates of the safety average time to dose reduction as indicated below:

### v2

@@
(June 17, 2025)
The majority of response-adaptive randomisation (RAR) designs in the literature rely on efficacy data to guide dynamic patient allocation. However, their applicability becomes limited in settings where efficacy outcomes, such as overall survival, are observed with a random delay. To address this limitation, we introduce SAFER (Safety-Aware Flexible Elastic Randomisation), a novel RAR design that leverages early-emerging safety data to inform treatment allocation decisions, particularly in oncology trials. The design is broadly applicable to contexts where prioritising the arm with a superior safety profile is desirable. This is especially relevant in non-inferiority trials, which aim to demonstrate that an experimental treatment is not inferior to the standard of care, while potentially offering improved safety or tolerability. In such trials, a well-established and unavoidable trade-off arises: maintaining statistical efficiency for the primary non-inferiority hypothesis while integrating safety-driven adaptations through RAR. The SAFER design addresses this trade-off by dynamically adjusting the allocation proportion based on the observed association between safety and efficacy endpoints. We illustrate the performance of SAFER through an extensive simulation study inspired by the CAPP-IT Phase III oncology trial. Results show that SAFER preserves statistical power when the safety-efficacy association is weak or absent and yields power gains when a strong positive association exists. Furthermore, SAFER reduces the adverse event rate during follow-up and offers flexible adaptation speed depending on the temporal alignment of the endpoints, faster when aligned and slower when misaligned.
Keywords FDA Project Optimus, Efficacy, Toxicity, Adverse events, Endpoints association, Patient-benefit
@@ 1 Introduction
In contemporary drug development, optimal dosing has become a central focus, aiming for a precise, patient-centred approach in oncology. Initiatives like the FDA’s Project Optimus [1] underscore this shift, emphasizing the best balance between drug’s efficacy and safety to maximize patient benefit and minimise toxicity. Traditionally, early-phase clinical trials primarily prioritise safety, often focusing on major adverse events, with efficacy being largely exploratory. As development progresses, later phases are primarily driven by efficacy decisions. Although comprehensive safety data are collected, this information is typically not formally integrated into design decisions, despite the goal of improving the overall benefit-risk profile and enabling patients to remain on treatment longer.
@@
This becomes particularly relevant in the context of the modern oncology treatment landscape, which has been transformed by the advent of targeted therapies, monoclonal antibodies, and immune checkpoint inhibitors. In fact, their administration typically spans prolonged treatment cycles and is often accompanied by chronic, low-grade adverse events (AEs), including fatigue, gastrointestinal disturbances, and dermatologic reactions. These AEs, although not classified as dose-limiting in the short windows of phase I trials, may significantly impair patient adherence, quality of life, and long-term tolerability in the later phases [2, 3]. Moreover, novel agents, designed to modulate specific molecular pathways, often exhibit plateauing efficacy at doses well below the maximum tolerated dose (MTD) [4, 5], underscoring the need for a revised perspective on the interplay between safety and efficacy outcomes. A pertinent example to illustrate this is sotorasib (LUMAKRAS), a KRAS G12C inhibitor approved for non-small cell lung cancer [6]. For this treatment a 960 mg dose was selected for pivotal trials. Subsequent analyses showed that a 240 mg dose offered similar efficacy with a better safety profile, leading the FDA to mandate a post-marketing dose optimization study.
@@
The above rationale motivates our interest in the following crucial question: how can a late phase adaptive design help identify alternative dosing strategies such as lower or fractionated doses, different schedules or regiments deliver, with comparable or superior efficacy and fewer side effects? This is particularly relevant to address the potential impact on phase II and III studies from dosing decisions from early trials often adopted on immediate and severe safety data. This is especially the case for agents with delayed efficacy signals and cumulative toxicity profiles that influence long-term adherence and outcomes.
@@
Our work aims to address the research question presented above by leveraging Response-Adaptive Randomisation (RAR). RAR refers to a class of statistical methodologies used in clinical trial design to dynamically adjust the probability of assigning patients to different treatment arms based on accumulating outcome data [7]. By increasingly favouring arms showing early superior performance, RAR can enhance both patient outcomes and the statistical efficiency of trials [8]. This adaptability is particularly advantageous when outcomes are available shortly after treatment initiation, allowing real-time modification of allocation proportions. However, in trials where primary efficacy outcomes are delayed, such as overall survival in oncology, RAR becomes more challenging due to the unavailability of timely, informative data during enrolment. Several strategies have been proposed to address this, including the use of early surrogate endpoints that correlate with long-term outcomes [9, 10, 11]. This paper introduces a RAR design that, alternatively, uses early-emerging safety data to inform treatment allocation decisions in oncology trials.
However, a critical component of a safety-driven RAR design is the consideration of the potential interplay between safety and efficacy endpoints. While some drugs may maintain or even improve efficacy after dose reductions, others could exhibit a diminished therapeutic benefit. From a statistical standpoint, RAR designs may introduce challenges in preserving statistical power, especially under certain conditions. Although equal allocation does not universally guarantee maximal power in superiority trials, as highlighted by recent work by Pin et al. (2024) [12], it remains an effective and commonly adopted strategy in many trial settings. On the contrary, in the context of non-inferiority trials with binary (Bernoulli) or time-to-event (exponential) outcomes, a balanced 1:1 allocation ratio is typically optimal due to the mathematical structure of the variance, which is a function of the mean. Under the alternative hypothesis in such trials, the non-inferiority assumption postulates that the average treatment effects across the intervention and control arms are clinically equivalent, within a narrowly defined margin of non-inferiority. This margin is usually small [13], and as a result, the outcome variances in both arms are expected to be similar. Given this, allocation strategies that aim to minimize the variance of the estimated treatment effect, such as Neyman allocation [14, 15], which assigns subjects in proportion to the standard deviation of the outcome in each arm, tend to approximate or even coincide with balanced allocation in these settings. Thus, attempts to skew allocation dynamically based on accumulating data, as done in RAR designs, may deviate from this optimal variance-reducing strategy. Consequently, an unavoidable and well-established trade-off arises for such designs: to balance the goals of preserving inferential efficiency for the primary non-inferiority outcome while incorporating safety considerations into the randomisation process through RAR.
@@
To this regards, Baldi Antognini et al. (2020) [16] introduced a compound optimal allocation strategy in survival endpoint trials to address the trade-off between "ethical" and inferential concerns in the choice of the randomisation proportion. Although their work is aligned with our rationale, it is limited to single-outcome efficacy trials since it does not capture the complexity of a joint safety-efficacy decisions.
@@
Therefore, we propose an innovative Safety-Aware Flexible Elastic Randomization (SAFER) design. SAFER dynamically adjusts patient allocation proportions based on accumulating safety endpoint data, with the aim of leveraging the association between efficacy and safety outcomes to balance clinical benefit and statistical validity. This adaptive strategy is particularly suitable for late-phase clinical trials, as it enhances patients’ treatment experience while preserving statistical power.
@@
The remainder of this paper is structured as follows. Section 2 introduces the key components of the SAFER design: (i) the efficacy endpoint, including the associated power function and sample size calculation; (ii) the definition of the safety endpoint; (iii) the target allocation proportion; and (iv) the integration of efficacy and safety data to guide the targeted allocation proportion. Section 3 introduces the CAPP-IT study as a motivating example and presents its re-design in a non-inferiority framework using the SAFER approach. Specifically, we outline the rationale and report the results from several distinct simulation scenarios, aimed at evaluating the performance of the proposed design under varying conditions. Section 4 discusses the simulation findings, explores broader implications of the SAFER design, and concludes with a summary of key contributions and directions for future research.
2 SAFER design components
2.1 Efficacy Endpoint
@@
Our proposed design is specifically tailored to accommodate for the typical primary endpoint used in Phase II/III oncology trials. Specifically, we consider a time-to-event measure, such as Progression-Free Survival (PFS). Consequently, both the sample size and the primary analysis of our design will be expressed throughout this paper in terms of this designated endpoint. The sample size will be determined accordingly to the expected effect size to ensure adequate statistical power for that main primary analysis. We will now formally outline this rationale, including the underlying assumptions and notation. The exponential distribution was assumed for simulation purposes; the implications and limitations of this assumption are further discussed in the Discussion section.
@@
Let $j=E,C$ denote the experimental treatment and control arms, respectively. The patients’ primary efficacy endpoint (e) i.e., their survival time ($Y^{(e)}$ ) is assumed to be independent and identically distributed (i.i.d.) with probability density function (p.d.f) when treatment $j$ is assigned:
@@
where $\theta_{j}^{(e)-1}=\lambda_{j}^{(e)}\in\mathbb{R}^{+}$ denotes the event rate. Therefore, $\theta_{j}^{(e)}$ represents the mean of the treatment effect, and $\theta_{j}^{(e)2}$ represents the variance of the treatment effect. We focus on the case in which the study is testing for the non-inferiority of the experimental with respect to the control arm. The null hypothesis for the mean difference can be expressed as [17, 18]
@@
$W=\frac{\hat{\theta}_{C}^{(e)}-\hat{\theta}_{E}^{(e)}-m}{\sqrt{Var(\hat{\theta}_{C}^{(e)}-\hat{\theta}_{E}^{(e)})}}\sim\mathcal{N}(0,1),$ (3)
@@
$\sqrt{Var(\hat{\theta}_{C}^{(e)}-\hat{\theta}_{E}^{(e)})}=\sqrt{N^{-1}\cdot(\hat{\theta}_{C}^{2}+\hat{\theta}_{E}^{2})},$ (4)
with $N$ equal to the sample size. Although both the semi-parametric Cox model and the parametric exponential survival models are consistent estimators under exponentiality, in practice, the semi-parametric model often converges more quickly and accurately in simulations [19, 20]. This is primarily due to its reliance on partial likelihood, fewer distributional assumptions, and better numerical conditioning in the presence of non-informative censoring and finite samples. Therefore, we estimate the Wald statistic through the Cox model [21]. We also incorporate an early stopping feature into our designs, recognizing that this feature, when combined with RAR, can offer advantages over RAR alone [22]. For simplicity, we have focused on only two analyses: an interim analysis (k=1) and the final analysis (k=2). Therefore, we will provide test statistics $W_{k=1}$ and $W_{k=2}$, estimated based on the first group available at the interim analysis and the complete sample at the end of follow-up, respectively.
To estimate the minimum number of observed progression events $P$ needed for the study, we construct a $H_{0}$ rejection region as $R=\left\{-\infty,z_{1-\beta}\right\}$, with
@@
where $\alpha$ is the significance level, $\beta$ is the type-II Error rate, $HR_{H_{1}}$ is the hazard ratio ($\lambda_{E}/\lambda_{C}$) under the alternative hypothesis, and $\pi_{E}$ is the allocation proportion to the experimental arm (E). In Section 3 we vary $\pi_{E}$ from 0.5 to 0.8. Therefore, given the rejection region, we can calculate the minimum number of events $P$ needed to test the non-inferiority on the primary endpoint between treatments, based on the following power function definition [23].
$Power=P_{H_{1}}(Z_{1}\in\mathcal{R}_{1})+P_{H_{1}}(Z_{1}\in\mathcal{C}_{1},Z_{2}\in\mathcal{R}_{2})$ (6)
In equation (6), $Z_{1}$ and $Z_{2}$ are the statistic tests at the interim and final analyses. $P_{H_{1}}(Z_{1}\in\mathcal{R}_{1})$ is $power_{1}$ (interim analysis power) i.e., the probability at the interim, under the alternative hypothesis ($H_{1}$), to fall into $R_{1}$, the region of rejection of the null hypothesis ($H_{0}$). Moreover, $P_{H_{1}}(Z_{1}\in\mathcal{C}_{1},Z_{2}\in\mathcal{R}_{2})$ is $power_{2}$ (final analysis power) i.e., the probability, under $H_{1}$ of reject $H_{0}$ at the final analysis given that the study continued at interim. Therefore, $power_{2}$ is given by a bivariate integral where $R_{2}$ and $C_{1}$ ($1-R_{1}$) indicate the rejection regions of $H_{0}$ and continuation at the final and interim analyses, respectively.
After the calculation of the minimum number of events $P$ to reach an overall power of at least 0.8 (assuming $\alpha$ = 0.05, $\pi_{E}$ = 0.5, the $H_{1}$ and the $m$ as defined below in Section 3), we adjust the total number of events needed due to the application of the Lan-DeMets O’Brien-Fleming $\alpha$-spending function [24] to account for $k=2$ analyses. Finally, due to the survival primary endpoint, the overall sample size ($N$) is estimated by dividing the total number of events needed ($P$) by the probability to observe at least one event during the follow-up, i.e., $Pr(\text{at least one event in time }t)=1-e^{-\lambda\cdot t}$ where lambda is the weighted (across arms) average event rate expected in the sample and t the maximum follow-up time.
@@ 2.2 Safety Endpoint
The safety endpoint in our design is the time to dose-reduction/drug discontinuation due to low drug tolerability or AEs. The patients’ safety endpoint (s) ($Y^{(s)}$), is assumed to be i.i.d., with p.d.f. when treatment $j$ is assigned:
@@
2.3 Allocation proportion
For our proposed design, the Neyman allocation [14, 15, 7] is a reasonable choice of allocation proportion to target. First, the Neyman allocation, maximizes the statistical power of a widely used Wald test to detect the true difference between arms; second, under the assumption of exponentially distributed survival endpoint, the Neyman proportion will consistently assign the majority of patients to the treatment that patients are observed to discontinue later [16] (or tolerate for the longest). Note that the latter is not always true for other endpoint distributions, such as the binary case [25], where the Neyman allocation can favour the worst treatment to minimize the variability of the Wald test. In our design, the Neyman target allocation proportion is estimated based on the estimated safety average time to dose reduction as indicated below: